an actor invest some money at 5% and 30000 more than twice the amount at 10% the total annual interest earned from the investment is 36000 how much did he invent at each amount
Answer by lwsshak3(11628)

an actor invest some money at 5% and 30000 more than twice the amount at 10% the total annual interest earned from the investment is 36000 how much did he invent at each amount.
***
let x=amt invested at 5%
2x+30000=amt invested at 10%
..
.05x+.10(2x+30000)=36000
.05x+.20x+3000=36000
0.25x=36000-3000
.25x=33000
x=132000
2x+30000=264000+30000=294000
..
amt invested at 5%=132000
amt invested at 10%=294000